# Towards Generalizable Autonomous Penetration Testing via Domain Randomization and Meta-Reinforcement Learning

### v1

@@ IV-B1 Model penetration testing as an MDP
$\mathcal{A}$ represents the set of all actions available to the agent. For pentesting tasks, these actions usually include various forms of information gathering, system probing, and vulnerability exploitation. In contrast to previous research that used abstract actions, our approach emphasizes the use of concrete and executable actions for end-to-end autonomous pentesting. These actions are aligned with specific commands or payloads derived from real pentesting tools or systems, such as Kali [^3] https://www.kali.org/, Metasploit [^4] https://www.metasploit.com/, Nmap [^5] https://nmap.org/, and so on. Note that a larger action space necessitates more extensive exploration for agents to discover effective actions, thereby potentially reducing the tractability of the agent’s training process. This phenomenon can be seen in experiments in Section V-D.
@@
IV-B2 Simulated environment construction
@@ IV-B3 Policy training
PPO is an on-policy actor-critic method with two primary variants of PPO: PPO-Penalty and PPO-Clip, where we focus on the most widely used variant PPO-Clip. By using PPO, the agent updates policy $\pi_{\phi}$ by taking multiple steps of (usually minibatch) SGD to minimize the following objective:
@@
IV-C LLM-powered Domain Randomization
@@ IV-D Meta-RL Training
V Experiments
V-A Experimental Setup
V-A1 Penetration testing environments
@@
Table I lists all experimental vulnerability environments sourced from Vulhub, each serving as an original training environment $\mathcal{M}^{k}_{0}$ in GAP, where $k$ represents the CVE identifier (CVE ID). We selected these vulnerability environments for training and testing the performance of the pentesting agent because they target various representative vulnerable products. And more importantly, we have reliable and stable exploits of these vulnerabilities available for the agent to use. Theoretically, our approach can be extended to a broader range of vulnerability environments.
V-A2 Agent settings
@@
In line with Equation 4, our reward function is formulated based on subjective estimates provided by human experts. Specifically, a reward of +1000 is assigned when the agent successfully exploits the correct vulnerability and compromises the target host. Additionally, a reward of +100 is allocated for proficiently gathering information about the target, such as the OS type, version, running services, and website fingerprint. Conversely, incorrect or illogical actions result in a penalty of -20.
V-A3 Implementation details
@@
We employ the classical PPO algorithm for policy learning and the MAML algorithm for meta-RL training. The synthetic environments used for meta-RL training are generated using GLM-4 [53] due to its satisfactory performance, and we assess their validity based on expert evaluators who are familiar with real-world pentesting scenarios. In fact, as a simple and general framework, GAP is not restricted to specific RL or meta-RL algorithms, nor is it limited to particular LLMs.
@@
TABLE I: Vulnerability environments.
CVE ID Vulnerability Name
CVE-2018-7600 Drupal Core Remote Code Execution Vulnerability
CVE-2019-9193 PostgreSQL Authenticated Remote Code Execution Vulnerability
CVE-2020-10199 Sonatype Nexus Repository Remote Code Execution Vulnerability
CVE-2020-7247 OpenSMTPD Remote Code Execution Vulnerability
CVE-2020-7961 Liferay Portal Deserialization of Untrusted Data Vulnerability
CVE-2021-22205 GitLab Unauthenticated Remote Code Execution Vulnerability
CVE-2021-25646 Apache Druid Remote Code Execution Vulnerability
CVE-2021-3129 Laravel Ignition File Upload Vulnerability
CVE-2022-0543 Debian-specific Redis Server Lua Sandbox Escape Vulnerability
CVE-2023-32315 Ignite Realtime Openfire Path Traversal Vulnerability
V-B Research Questions
GAP aims to improve the agent’s generalization ability, thereby allowing the agent to draw inferences from one to another. We aim to answer the following research questions (RQs) from our experiments and analysis:
@@ V-B Research Questions
V-C Evaluation Metrics
V-C1 Evaluation of RQ1
@@
To investigate RQ1, in each run, we evaluate the agent’s learning curve and training time for all original training environments as listed in Table I. We compare the performance of the agent with different sizes of action spaces ($\left|\mathcal{A}\right|\in{\left\{100,500,1000\right\}}$). We implement three independent runs using different seeds, with average results shown in Section V-D. Note that in subsequent experiments, the agent’s action space is fixed to 1000.
V-C2 Evaluation of RQ2
@@
To investigate RQ2, following [5, 48], we assess the agent’s zero-shot generalization performance in both training and testing environments, which measures its ability to bridge the generalization gap (calculated by Equation 3 ). Besides, the average success rate is also used as a metric, which reflects the proportion of agent successfully compromised hosts to the total number of hosts in testing environments.
@@
To this end, we create three variants for each $\mathcal{M}^{k}_{0}$ in Table I by changing its host configurations to serve as testing environments denoted by $\mathcal{M}^{k}_{test}=\left\{\mathcal{M}^{k}_{1},\mathcal{M}^{k}_{2},\mathcal{M}^{k}_{3}\right\}$. These variants in $\mathcal{M}^{k}_{test}$ are similar to $\mathcal{M}^{k}_{0}$ as they share the same vulnerability $k$, but they differ in host configurations such as ports, services, operating system versions, website fingerprints, and others. These differences create reality gaps that mimic the diversity and variability found in real-world host environments.
@@
In each implementation run, we train the agent for each $\mathcal{M}^{k}_{0}$ in Table I , and later evaluate its zero-shot performance in $\mathcal{M}^{k}_{0}$ and testing environments $\mathcal{M}^{k}_{test}$, as well as their pentesting success rate in $\mathcal{M}^{k}_{test}$. We implement three independent runs using different seeds, with average results shown in Section V-E.
V-C3 Evaluation of RQ3
@@
A well-generalized agent should also be able to effectively improve its learned policy to adapt to a new environment. To investigate RQ3, following [5, 49], we evaluate the agent’s few-shot policy adaption performance in dissimilar environments by allowing the agent to interact with a new environment online.
@@
In particular, in each run, for each $\mathcal{M}^{k}_{0}$ in Table I, we randomly select another environment $\mathcal{M}^{k^{\prime}}_{0}$ as the testing environment, where $k^{\prime}\neq k$. The agent is trained in $\mathcal{M}^{k}_{0}$ and fine-tuned in $\mathcal{M}^{k^{\prime}}_{0}$. We evaluate its learning curve and training time in the testing environment, which measures how quickly the agent can improve its performance on a new task.
@@
In addition, following [54], we also evaluate the agent’s jumpstart performance, which refers to the agent’s initial performance when it starts interacting with the testing environment. It assesses how quickly the agent can abstract broader knowledge from past experiences and apply it to novel situations, thereby adapting its policy and achieving meaningful rewards in the early stages of training. We implement three independent runs using different seeds, with average results shown in Section V-F.
V-D Policy Learning in Real Environments (RQ1)
@@
Being able to learn in an unknown environment is the foundation skill for pentesting agents. In this part, we train the agent to learn policies from scratch in all original training environments that are unknown to the agent. We compare the learning performance of the agent across different action spaces ($\left|\mathcal{A}\right|\in{\left\{100,500,1000\right\}}$). Fig.6 and Fig.7 display the agent’s learning curve and training time, respectively.
@@
Figure 7: Training time of the agent with varying action space sizes in all original training environments over three independent runs. The box plot summarizes the distribution of training time across 500 episodes for each action space. The dashed line connects the average training time across different action spaces.
The experimental results demonstrate that the agent can learn policy in unknown environments. However, as the action space increases, there is a decrease in learning efficiency, a slower convergence rate of the learning curve, and an increase in training time accordingly. This phenomenon arises because a larger action space requires more extensive exploration to identify effective actions, thereby prolonging the time needed for the agent to discover the optimal policy. Additionally, optimizing the policy model becomes more computationally demanding with increased iterations and computational effort as the action space expands.
@@
V-E Zero-shot Policy Transfer in Similar Environments (RQ2)
In this part, we demonstrate the agent’s zero-shot generalization ability in testing environments similar to the original training environment.
In particular, the agent is pre-trained using various methods in the original training environment and subsequently transferred its policies to its training environment and testing environments. We compare the performance differences between GAP and the baseline method, where the baseline method represents directly evaluating the agent’s performance in testing environments. Additionally, we also analyze the effect of the number of meta-training environments on the zero-shot generalization performance of GAP. We use $n$-GAP to denote the use of $n$ meta-training environments in the GAP, where $n\in\left\{3,5,8\right\}$ in our experiments.
Figure 8: Zero-shot generalization performance in original training environments and testing environments. The solid line represents the agent’s average success rate in testing environments.
As results show in Fig.8, we can find that the agent trained using the baseline method exhibits a lower pentesting success rate in testing environments, accompanied by noticeable performance loss, indicating a generalization gap. In contrast, our proposed framework, GAP, significantly enhances success rates in testing environments compared to the baseline method and bridges the generalization gap, demonstrating improved zero-shot generalization ability.
@@ V-E Zero-shot Policy Transfer in Similar Environments (RQ2)
V-F Few-shot Policy Adaption in Dissimilar Environments (RQ3)
In this part, we evaluate the agent’s policy adaption ability when faced with testing environments that are dissimilar to their original training environments. The policy adaption ability refers to how well the agent improves its performance through fine-tuning or adjusting its existing trained model in a new environment.
We use the following three approaches to train or fine-tune the agent with action space $\left|A\right|=1000$ in new environments:
  • Learn from scratch. The agent is trained in the new environment without any prior knowledge.
  • Fine-tune baseline. We pre-train the agent in the original training environment and directly fine-tune it in the new environment.
  • Fine-tune GAP. Following the Real-to-Sim-to-Real pipeline, the agent is pre-trained in the original training environment, and then its generalization is improved by adopting domain randomization and meta-RL training. Finally, we fine-tune it in the new environment.
Figure 9: Learning curves of the agent in testing environments over three independent runs. Each curve is the average reward, while the shaded area denotes the 95% confidence intervals.
@@ V-F Few-shot Policy Adaption in Dissimilar Environments (RQ3)
Fig.9 and Fig.10 display the agent’s learning curve and training time in new environments, respectively. The results show that compared to learning from scratch, fine-tuning baseline and fine-tuning GAP demonstrate significantly improved policy adaption ability in new environments. This is mainly evidenced by their quicker convergence in learning curves and reduced training time. Among the three methods studied, fine-tuning GAP stands out as the best, notably reducing average training time by approximately 38% compared to learning from scratch (from 4045 seconds down to 2500 seconds).
In particular, from Fig.9 we can see that GAP demonstrates superior jumpstart performance compared to the baseline method, enabling the agent to explore effective actions more rapidly during initial training in new environments. This ability suggests that the agent can generalize broader knowledge from past experiences and apply it effectively to novel situations. This is attributed to the introduction of a meta-learning mechanism, empowering the agent to enhance its capability for policy adaptation by learning how to learn.
VI Discussion

### v2

@@ 4.2.1 Model penetration testing as an MDP
$\mathcal{A}$ represents the set of all actions available to the agent. For pentesting tasks, these actions usually include various forms of information gathering, system probing, and vulnerability exploitation. In contrast to previous research that used abstract actions, our approach emphasizes the use of concrete and executable actions for end-to-end autonomous pentesting. These actions are aligned with specific commands or payloads derived from real pentesting tools or systems, such as Kali [^3] https://www.kali.org/, Metasploit [^4] https://www.metasploit.com/, Nmap [^5] https://nmap.org/, and so on. Note that a larger action space necessitates more extensive exploration for agents to discover effective actions, thereby potentially reducing the tractability of the agent’s training process. This phenomenon can be seen in experiments in Section 6.1.
@@
4.2.2 Simulated environment construction
@@ 4.2.3 Policy training
PPO is an on-policy actor-critic method with two primary variants of PPO: PPO-Penalty and PPO-Clip, where we focus on the most widely used variant PPO-Clip. By using PPO, the agent updates policy $\pi_{\phi}$ by taking multiple steps of (usually minibatch) SGD to optimize the following objective:
@@
4.3 LLM-powered Domain Randomization
@@ 4.4 Meta-RL Training
5 Experimental Settings
5.1 Penetration testing environments
@@
Table 1 lists all experimental vulnerability environments sourced from Vulhub, each serving as an original training environment $\mathcal{M}^{k}_{0}$ in GAP, where $k$ represents the CVE identifier (CVE ID). We selected these vulnerability environments for training and testing the performance of the pentesting agent because they target various representative vulnerable products. And more importantly, we have reliable and stable exploits of these vulnerabilities available for the agent to use. Theoretically, GAP can be extended to a broader range of vulnerability environments.
5.2 Agent settings
@@
Following [13], our reward function is formulated based on subjective estimates provided by human experts. Specifically, a reward of +1000 is assigned when the agent successfully exploits the correct vulnerability and compromises the target host. Additionally, a reward of +100 is allocated for proficiently gathering information about the target, such as the OS type, version, running services, and website fingerprint. Conversely, incorrect or illogical actions result in a penalty of -10.
5.3 Implementation details
@@
In GAP, we employ the classical PPO algorithm for policy learning and the MAML algorithm for meta-RL training. The synthetic environments used for meta-RL training are generated using GLM-4 [56] due to its satisfactory performance, and we assess their validity based on expert evaluators who are familiar with real-world pentesting scenarios. In fact, as a simple and general framework, GAP is not restricted to specific RL or meta-RL algorithms, nor is it limited to particular LLMs.
@@
Table 1: Vulnerability environments.
CVE ID Vulnerability Name Vulnerable Product
CVE-2018-7600 Drupal Core Remote Code Execution Vulnerability Drupal Core
CVE-2019-0230 Apache Struts OGNL Remote Code Execution Vulnerability Apache Struts
CVE-2019-9193 PostgreSQL Authenticated Remote Code Execution Vulnerability PostgreSQL
CVE-2020-10199 Sonatype Nexus Repository Remote Code Execution Vulnerability Nexus
CVE-2020-7247 OpenSMTPD Remote Code Execution Vulnerability OpenSMTPD
CVE-2020-7961 Liferay Portal Deserialization of Untrusted Data Vulnerability Liferay Portal
CVE-2020-9496 Apache OFBiz Remote Code Execution Vulnerability Apache OFBiz
CVE-2020-16846 SaltStack Salt Shell Injection Vulnerability SaltStack Salt
CVE-2021-22205 GitLab Unauthenticated Remote Code Execution Vulnerability GitLab
CVE-2021-25646 Apache Druid Remote Code Execution Vulnerability Apache Druid
CVE-2021-3129 Laravel Ignition File Upload Vulnerability Ignition
CVE-2021-41773 Apache HTTP Server Path Traversal Vulnerability Apache HTTP Server
CVE-2022-0543 Debian-specific Redis Server Lua Sandbox Escape Vulnerability Redis
CVE-2022-22947 VMware Spring Cloud Gateway Code Injection Vulnerability Spring Cloud Gateway
CVE-2022-46169 Cacti Command Injection Vulnerability Cacti
CVE-2023-21839 Oracle WebLogic Server Unspecified Vulnerability Oracle WebLogic
CVE-2023-32315 Ignite Realtime Openfire Path Traversal Vulnerability Openfire
CVE-2023-38646 Metabase Pre-Auth Remote Code Execution Vulnerability Metabase
CVE-2023-42793 JetBrains TeamCity Authentication Bypass Vulnerability JetBrains
CVE-2023-46604 Apache ActiveMQ Deserialization of Untrusted Data Vulnerability Apache ActiveMQ
5.4 Research Questions
GAP aims to improve the agent’s generalization ability, thereby allowing the agent to draw inferences about other cases from one instance. We aim to answer the following research questions (RQs) from experiments and analysis:
  • 1.
@@ 5.4 Research Questions
5.5 Evaluation Metrics
Table 2 presents the training and testing environments, along with the metrics used to investigate all research questions. The details are as follows.
5.5.1 Evaluation of RQ1
@@
To investigate RQ1, in each run, we evaluate the agent’s learning curve and training time for all original training environments as listed in Table 1. We compare the performance of the agent with different sizes of action spaces ($\left|\mathcal{A}\right|\in{\left\{100,500,1000\right\}}$). We implement three independent runs using different seeds for each vulnerability environment, with average results shown in Section 6.1.
5.5.2 Evaluation of RQ2
@@
To investigate RQ2, following [16, 51], we assess the agent’s zero-shot generalization performance in both training and testing environments, which measures its ability to bridge the generalization gap (GenGap, calculated by Equation 3 ). Besides, the average success rate is also used as a metric, which reflects the proportion of agent successfully compromised hosts to the total number of hosts in testing environments.
@@
To this end, we create three variants for each $\mathcal{M}^{k}_{0}$ in Table 1 by changing its host configurations to serve as testing environments denoted by $\mathcal{M}^{k}_{test}=\left\{\mathcal{M}^{k}_{1},\mathcal{M}^{k}_{2},\mathcal{M}^{k}_{3}\right\}$. These variants in $\mathcal{M}^{k}_{test}$ are similar to $\mathcal{M}^{k}_{0}$ as they share the same vulnerability $k$, but they differ in host configurations such as ports, services, operating system versions, website fingerprints, and others. These differences create reality gaps that mimic the diversity and variability found in real-world host environments.
@@
In each implementation run, we train the agent for each $\mathcal{M}^{k}_{0}$ in Table 1 , and later evaluate its zero-shot performance in $\mathcal{M}^{k}_{0}$ and testing environments $\mathcal{M}^{k}_{test}$, as well as their pentesting success rate in $\mathcal{M}^{k}_{test}$. We implement three independent runs using different seeds for each vulnerability environment, with average results shown in Section 6.2.
5.5.3 Evaluation of RQ3
@@
A well-generalized agent should also be able to effectively improve its learned policy to adapt to a new environment. To investigate RQ3, following [16, 52], we evaluate the agent’s few-shot policy adaptation performance in dissimilar environments by allowing the agent to interact with a new environment online.
@@
In particular, in each run, for each $\mathcal{M}^{k}_{0}$ in Table 1, we randomly select another environment $\mathcal{M}^{k^{\prime}}_{0}$ as the testing environment, where $k^{\prime}\neq k$. The agent is trained in $\mathcal{M}^{k}_{0}$ and transfers its learned policy in $\mathcal{M}^{k^{\prime}}_{0}$. We evaluate its learning curve and training time in the testing environment, which measures how quickly the agent can adapt its performance to a new task.
@@
In addition, following [57], we also evaluate the agent’s jumpstart performance, which refers to the agent’s initial performance when it starts interacting with the testing environment. It assesses how quickly the agent can abstract broader knowledge from past experiences and apply it to novel situations, thereby adapting its policy and achieving meaningful rewards in the early stages of training. We implement three independent runs using different seeds for each vulnerability environment, with average results shown in Section 6.3.
Table 2: Training and testing environments and metrics for investigating research questions.
RQ Training Testing Metrics
RQ1 $\mathcal{M}^{k}_{0}$ Learning curve, Training time
RQ2 $\mathcal{M}^{k}_{0}$ $\mathcal{M}^{k}_{test}=\left\{\mathcal{M}^{k}_{1},\mathcal{M}^{k}_{2},\mathcal{M}^{k}_{3}\right\}$ GenGap, Success Rate
RQ3 $\mathcal{M}^{k}_{0}$ $\mathcal{M}^{k^{\prime}}_{0},k^{\prime}\neq k$ Learning curve, Training time, Jumpstart
6 Experimental Results
6.1 Policy Learning in Real Environments (RQ1)
@@
Being able to learn in various real environments is the foundation skill for pentesting agents. In this part, we train the agent to learn policies from scratch in all original training environments that are unknown to the agent. We compare the learning performance of the agent across different action spaces ($\left|\mathcal{A}\right|\in{\left\{100,500,1000\right\}}$). Fig.6 and Fig.7 display the agent’s learning curve and training time, respectively.
@@
The experimental results demonstrate that the agent can learn policy in training environments. However, as the action space increases, there is a decrease in learning efficiency, a slower convergence rate of the learning curve, and an increase in training time accordingly. This phenomenon arises because a larger action space requires more extensive exploration to identify effective actions, thereby prolonging the time needed for the agent to discover the optimal policy. Additionally, optimizing the policy model becomes more computationally demanding with increased iterations and computational effort as the action space expands.
@@
During interaction with the original training environments, the agent can gather the host configuration data (e.g., ports, services, web fingerprint, etc.) and save them in JSON format as the simulated copy of the real environment. This approach not only provides realistic simulated environments for agent training, effectively addressing training environment dilemma, but also enables the synthesis of additional diverse simulation environments using large language models, thereby augmenting the environment variety. An example of this process is shown in Fig.5, and more details of the simulated environments are available in our code repository.
Figure 7: Training time of the agent with varying action space sizes in all original training environments. The box plot summarizes the distribution of training time across 500 episodes for each action space. The dashed line connects the average training time across different action spaces.
6.2 Zero-shot Policy Transfer in Similar Environments (RQ2)
In this part, we demonstrate the agent’s zero-shot generalization ability in testing environments that have the same vulnerability as the original training environment.
In particular, the agent is pre-trained using various methods in the original training environment and subsequently transferred its policies to its training environment and testing environments. We compare the zero-shot generalization performance between GAP and two baseline methods: PPO [35] and April [13]. Both baselines are state-of-the-art methods that have been wildly used in autonomous pentesting.
Additionally, we also analyze the effect of the number of meta-training environments on the zero-shot generalization performance of GAP. We use $n$-GAP to denote the use of $n$ meta-training environments in the GAP, where $n\in\left\{3,5,8\right\}$ in our experiments.
Figure 8: Zero-shot generalization performance in original training environments and testing environments.
As shown in Fig.8, when the agent’s policy is transferred to the original training environments, all methods maintain satisfactory performance. However, when the policy is transferred to the testing environments, both baselines experience a noticeable performance loss. This performance gap between the training and testing environments is referred to as the generalization gap. Table 3 presents the generalization gap (GenGap) of different methods, along with their success rate in the testing environments. From the results we can see that our proposed framework, GAP, significantly enhances success rates in testing environments compared to the baseline method and bridges the generalization gap, demonstrating improved zero-shot generalization ability.
Table 3: Generalization gap and success rate in testing environments.
Method GenGap Success Rate
April 990.21 0.14
PPO 624.33 0.44
3-GAP 171.68 0.83
5-GAP 51.61 0.92
8-GAP 61.11 0.91
@@ 6.2 Zero-shot Policy Transfer in Similar Environments (RQ2)
6.3 Few-shot Policy Adaptation in Dissimilar Environments (RQ3)
In this part, we evaluate the agent’s policy adaptation ability when faced with testing environments that are dissimilar to the original training environments. The policy adaptation ability refers to the agent’s ability to quickly adapt its learned policy to unseen environments through policy transfer.
We use the following four methods to train the agent with action space $\left|A\right|=1000$:
  • 1.
    Learn from scratch. Similar to Section 6.1, the agent is directly trained using PPO in testing environments without policy transfer.
  • 2.
    PPO-Transfer. We pre-train the agent using PPO in original training environments and then transfer its learned policy to testing environments.
  • 3.
    April-Transfer. We pre-train the agent using April in original training environments and then transfer its learned policy to testing environments. April is the state-of-the-art autonomous pentesting framework that has been proven to exhibit strong transfer learning performance [13].
  • 4.
    GAP-Transfer. Following the Real-to-Sim-to-Real pipeline, the agent is pre-trained in the original training environment, and then its generalization is improved by adopting domain randomization and meta-RL training. Finally, we transfer its learned policy to testing environments.
Figure 9: Learning curves of different methods in testing environments.
Table 4: Jumpstart performance of different methods.
Method Jumpstart
GAP-Transfer 589.22
April-Transfer 367.21
PPO-Transfer 129.83
Learn from scratch 47.33
Fig.9 and Fig.10 display the agent’s learning curve and training time in testing environments, respectively. The results show that, compared to learning from scratch, the other three methods enhance convergence speed and reduce training time through policy transfer. Among them, GAP stands out as the most effective, reducing average training time by approximately 40% (from 4045 seconds to 2429 seconds).
Furthermore, as shown in Fig.9 and Fig.10 , it can be observed that while GAP-Transfer and April-Transfer exhibit similar convergence speeds, GAP-Transfer reduces training time by approximately 22% compared to April-Transfer (from 3129 seconds to 2429 seconds). This reduction is related to the jumpstart performance, as evidenced by Table 4 and Fig.11.
Table 4 presents the jumpstart performance of different methods, from which we can see that GAP-Transfer demonstrates superior jumpstart performance than others. Besides, Fig.11 illustrates a strong negative linear correlation between jumpstart performance and average training time, with a Pearson correlation coefficient (Pearson’s r) close to $-0.99$. A better jumpstart allows the agent to explore effective actions more quickly during initial training in new environments, indicating that the agent can generalize knowledge from past experiences and apply it effectively to new situations. The outstanding performance of GAP-Transfer can be attributed to the incorporation of a meta-learning mechanism, which empowers the agent to improve its policy adaptation ability by learning how to learn.
@@ 6.3 Few-shot Policy Adaptation in Dissimilar Environments (RQ3)
Figure 11: Correlation between jumpstart and average training time of different methods.
7 Discussion